VEGAS Team Meeting
Attendees: Richard, Anish, Mark, Joe, Ray, Jason, Melinda, Amanda, Dan, Paul, John
- ADC Calibration
- Mixed Mode switching Summary and Priority
- Release of 1500 MHz etc bandwidth config_tool patch
- Review current issues: CICADAGreenBankSpectrometerCurrentIssues
- Data Directory Cleanup
- ADC Calibration: Hong is still working out some final details; Jason will check on status after the meeting. Use of Bank H will now need to be coordinated around KFPA flexibly-scheduled observations.
- Mixed mode switching: Joe has a proposed solution which falls into two parts: a) doing the tracking in the presence of no packet errors, and b) working around dropped packets. This is "quite high" priority. Mark will review the effort required, and propose a schedule at the next meeting. Regardless, we need Bank A to be set up the same as the (single) mode specified for other banks; Melinda will release this at the next available opportunity. This may (will?) run into the "bank change" issue, but Ray is looking into that.
- 1500MHz config_tool change: will be released the the "Bank A configuration" change.
- See CICADAGreenBankSpectrometerCurrentIssues
- Amanda proposes deleting data taken before 1st October. John F. will check with the galactic center pulsar people to make sure their data is archived. The JUNK directory can simply be deleted.
- Please make sure to properly code your timesheets for the current time period and going forward.
- Software engineers should be using a GBT ops software CM&E account
- Electronic engineers should be using a GBT ops electronics CM&E account
- Commissioning activities should be getting charged to hardware support and testing (432510)
- scientists assisting with shared-risk commissioning should be charging an observer support account.
432540 for the most part. Accounts 142023 (GB-based staff) and 142013 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)